OpenStudy (anonymous):

What are the solutions? 16x^2 + 9 = 0

OpenStudy (helder_edwin):

first add -9 to both sides then divide both sides by 12

OpenStudy (anonymous):

\[16x^{2} + 9 = 0 \] \[16x^{2} =-9\] \[x^{2} =\frac{-9}{16}\] \[\sqrt{x^{2}} =\sqrt{\frac{-9}{16}}\] \[x =\frac{\sqrt{-9}}{\sqrt{16}}\] \[x =\frac{3i}{4}\]

OpenStudy (anonymous):

do you understand it?

OpenStudy (anonymous):

yea i do now

OpenStudy (anonymous):

if i give you one to do will you be able to do it?